# Plurals: A System for Guiding LLMs Via Simulated Social Ensembles

### v5

@@ 7. Ethical Considerations
The use of personas in research and design raises ethical concerns around misrepresentation and stereotyping (sun_building_2024, ). Ultimately, almost any technical representation of human behavior is “lossy” in some way. However, we tried to reduce homogenization by encouraging intersectional persona generation (Case Study 1). Nonetheless, the potential for misrepresentation is a valid concern. We frame these concerns as pro tanto harms—harms that “have some bearing on what we ought to do but that can be outweighed” (askell_ai_2020, ). As Askell writes, most systems have some non-zero harm (askell_ai_2020, ). So, we also need to consider what would be the alternative if that system did not exist. Imperfect representation should be weighed against that perspective not being considered at all.
@@ 8. Discussion
Plurals provides both a computing paradigm and a concrete, usable system for creating pluralistic artificial intelligence. By embracing a diversity of perspectives rather than seeking an illusory “view from nowhere,” Plurals highlights the potential for more pluralistic artificial intelligence systems. The core principle is what we term “interactional pluralism”. This is a pluralism that exists not only in the distribution of agent properties, but also in the protocols that govern their interactions. This is a fundamentally different kind of AI pluralism than existing typologies (sorensen_roadmap_2024, ).
@@
8.2. Plurals encourages reflective development.
@@
When an end-user creates a Plurals deliberation, they intentionally decide the parameters of the deliberation—such as who is in the deliberation and how Agents should deliberate. In this sense, Plurals encourages AI developers to consciously think about the audience that they are building for. This encourages more reflective development practices (dourish_reflective_2004, ).
8.3. Plurals opens new opportunities for studying the impact of AI on humans.
@@
Our system can contribute to a line of research on how exposure to AI ideas might impact humans (ashkinaze_how_2024, ; argyle_leveraging_2023, ). Specifically: (1) Under what conditions do simulated perspectives help humans make better decisions or come up with better ideas? and (2) Through what mechanisms do simulated perspectives influence people?
Human-centric use cases of Plurals can be input-focused or output-focused, mirroring uses of human deliberative mini-publics (warren_deliberative_1996, ; bachtiger_deliberative_2018, ; smith_mini-publics_2018, ). Input-focused applications use Plurals to inform human decisions, similar to Oregon’s mini-publics that provide voters with pros and cons lists (knobloch_did_2013, ). Output-focused applications treat Plurals output as the terminal endpoint.
  • Input-focused applications use Plurals deliberations as an input to inform humans. Examples: Brainstorming product design choices, policy options, multi-perspective revisions, and hypothesis generation. The key research questions here are around when and how such AI-generated inputs lead to better human decisions.
  • Output-focused applications focus on Plurals deliberations as the end-product. Examples: automated content generation, classification, and steerable moderation. Here, the questions are more around optimizing the quality and usefulness of the outputs, themselves.
@@
  • Plurals allows exploration of multi-LLM information diffusion dynamics (zhang_dynamics_2016, ; bakshy_role_2012, ).
@@
Our “interactional pluralism” can integrate with various AI alignment techniques. One integration we are particularly interested in is combining our approach with retrieval-augmented generation (RAG) and case-based reasoning (feng_case_2023, ; salemi_lamp_2024, ) to enable Agents to deliberate from diverse informational starting points, more closely approximating human deliberation. Also, future work could involve training models on multi-turn deliberations from different Structures and combination instructions, allowing models to more permanently “learn” from deliberative experiences. Finally, as interest in model abstentions (wen_know_2024, ) grows, to what extent can Plurals deliberations be used as steerable guardrails?
@@ 9. Conclusion
Future work on Plurals could explore a range of directions, such as: incorporating RAG into deliberation so Agents have distinct knowledge, using Plurals deliberations as moderation endpoints, using other data-based persona generation methods (li_steerability_2024, ), and conducting field studies to evaluate the impact of Plurals output in real-world settings. Broadly, we see the human-centric applications of Plurals as divided between serving as inputs for human decision-makers or creating outputs that are more helpful or resonant than standard methods.
@@
Appendix A Deliberation Ideals
@@
Table 2. Translating ideals of deliberative democracy into instructions for LLMs. Starting from the taxonomy in Bächtiger et al. (bachtiger_deliberative_2018, ), two authors engaged in an iterative process where we first screened ideals for relevance to AI agents and then translated ideals into LLM instructions.
@@
Appendix B Creating Custom Structures
Appendix C Case Study: Diversity of ANES Persona Responses
Political Issues
Generation
We prompted GPT-4o and Claude Sonnet to provide 100-word stances on each issue, varying ideology (liberal or conservative) and agent type (non-Plurals minimal prompt or Plurals ANES integration). For non-Plurals, we used the system instruction “You are a [liberal/conservative]”. For Plurals, we generated unique personas using our “ideology” initializer and “anes’ persona template (which tells the model how to enact this persona). Hence, the Plurals personas will have additional demographic information whereas the standard, non-Plurals persona only has ideology. We generated 30 responses for each (issue, ideology, agent type, model) combination.
Measures
We pooled the responses for each (issue, ideology, agent type, model) combination into a corpus and then represented this corpus as a bag of words, similar to (padmakumar_does_2024, ). We then measured the lexical diversity of Plurals vs Non-Plurals corpora. Intuitively, diverse responses would mean low repetition. The type-token ratio (TTR) (johnson_studies_1944, ) is a common measure of linguistic diversity. It is the number of unique tokens divided by the number of total tokens. When this ratio is high, words are relatively unique, and vice versa. We follow (padmakumar_does_2024, ) and compute this metric for various degrees of n-grams (1-grams, 2-grams, 3-grams, 4-grams, 5-grams). We also compute HD-D, which is a modification of TTR that adjusts for texts of varying lengths (mccarthy_vocd_2007, ).
Results
In an initial analysis, Plurals ANES responses had higher lexical diversity in 76 of 80 comparisons111111(2 ideologies x 4 issues x 5 n-gram orders x 2 metrics) for GPT-4o and all 80 comparisons for Claude Sonnet (Appendix Figure 5). These proportions (95% and 100%) significantly differ from chance (two-tailed exact binomial test, p ¡ .001). To account for correlations among diversity metrics, we conducted a secondary analysis using the first principal component from the 10 diversity metrics, which explained 88% of variance. A two-tailed permutation test on the difference in means for this component—aggregated at the (issue, ideology, agent type, model) level—rejected the null hypothesis at p ¡ .001. The mean paired difference (Plurals PC1 - Non-Plurals PC1) was $M=3.67,\text{95\% bootstrap CI}=[2.78,4.68],d_{z}=1.84$. These results confirm that augmenting prompts with demographic variables increases response diversity compared to ideological prompts alone.
@@ Results
Appendix D Case Study: Deliberation Instructions
D.1. Example Ballot Prompt
Argue for or against this ballot initiative.
DESCRIPTION
Prohibit carbon tax credit trading and repeal provisions of the 2021 Washington Climate Commitment Act (CCA),
a state law that provided for a cap and invest program designed to reduce greenhouse gas (GHG) emissions by 95% by 2050
VOTING
-A "yes" vote supports prohibiting any state agencies from implementing a cap and trade or cap and tax program
and repealing the 2021 Washington Climate Commitment Act (CCA), a state law that provided for a cap and invest
program designed to reduce greenhouse gas (GHG) emissions by 95% by 2050.
-A "no" vote opposes prohibiting state agencies from implementing a cap and trade or cap and tax program and opposes
repealing the 2021 Washington Climate Commitment Act (CCA), a state law that provided for a cap and invest
program designed to reduce greenhouse gas (GHG) emissions by 95% by 2050.
DETAILED OVERVIEW
[omitting for space]
Constraints
Answer in 150 words.
D.2. Combination Instructions
These were the combination instructions given to Agents.
D.2.1. Emotional
KEEP TRACK OF DEBATE HISTORY
You are in a debate with another agent. Here is what you have said and what the other agent has
said. Never refer to yourself in the third person.
<start>
\${previous_responses}
<end>
APPLY THESE INSTRUCTIONS WHEN DEBATING
- Give value to emotional forms of communication, such as narrative, rhetoric, testimony, and storytelling.
D.2.2. Rational Condition
KEEP TRACK OF DEBATE HISTORY
You are in a debate with another agent. Here is what you have said and what the other agent has
said. Never refer to yourself in the third person.
<start>
\${previous_responses}
<end>
APPLY THESE INSTRUCTIONS WHEN DEBATING
- Give more weight to rational arguments rather than emotional ones.
D.3. Task Wording
Below are excerpts from debates generated by two different AI systems. One AI system generated responses after being instructed: “Give more weight to rational arguments rather than emotional ones.” Another system generated responses after being instructed: “Give value to emotional forms of communication, such as narrative, rhetoric, testimony, and storytelling.” Please select which response most adheres to: “Give more weight to rational arguments rather than emotional ones.”
Appendix E Multilevel Logistic Regressions of Efficacy Studies
@@
Appendix F Case Study: Solar Panels
F.1. Commitment Check
@@
F.2. Plurals Code

### v6

@@ 7. Ethical Considerations
Ethical Arguments for Pluralistic AI
We acknowledge the ethical considerations that Plurals introduces and argue that pluralistic AI systems are ethically preferable to those that collapse diverse viewpoints into a single perspective (“output collapse”). Our system promotes accountability by requiring developers to explicitly specify Agent characteristics (Raji et al., 2020), enables upweighting minority voices through Structure connectivity, and demonstrates proof-of-concept capabilities through empirical studies. We show Plurals can reduce output homogenization (Study 1), implement steerable deliberation protocols (Study 2), generate output resonating with distinct audiences (Studies 3-5), and possibly support customizable moderation (Study 6). By giving users control over whose voices to include and how they interact, Plurals represents a meaningful step towards pluralistic AI systems. Nonetheless, we discuss some ethical considerations below.
@@
Risk of False Empathy
Recent design critiques argue that empathy-facilitating simulations that try to capture “being like” a target group are problematic (Bennett and Rosner, 2019). These simulations can: deny the authority of lived experience, create divisions between designers and users, and treat the simulated group as a spectacle. Like many simulations, Plurals’ use of personas carries these risks. But our emphasis is on deliberative exchanges between Agents rather than static snapshots. The deliberative nature of our system already acknowledges that simulated perspectives are necessarily partial.
Training Data Inequities
Training data constrains any AI system, including Plurals. When training data contains societal biases, LLMs risk reproducing these biases (Navigli et al., 2023). One approach we encourage, as we did in this paper, is to recruit actual members of a group to verify that representations are resonant with that group. Moreover, LLMs may be worse at modeling groups who appear rarely in training data (Kandpal et al., 2023), echoing “design exclusion” (Keates and Clarkson, 2002; Clarkson et al., 2003) concerns in HCI. The representational gap may be partially reduced through steering that Plurals affords (section 6) or by selecting LLMs with more ethical/transparent training data practices.
@@
The use of personas in research and design raises ethical concerns around misrepresentation and stereotyping (Sun et al., 2024). Ultimately, almost any technical representation of human behavior is “lossy” in some way. However, we tried to reduce homogenization by encouraging intersectional persona generation (Case Study 1). Nonetheless, the potential for misrepresentation is a valid concern. We frame these concerns as pro tanto harms—harms that “have some bearing on what we ought to do but that can be outweighed” (Askell, 2020). As Askell writes, most systems have some non-zero harm (Askell, 2020). So, we also need to consider what would be the alternative if that system did not exist. Imperfect representation should be weighed against that perspective not being considered at all.
@@ 8. Discussion
Plurals provides both a computing paradigm and a concrete, usable system for creating pluralistic artificial intelligence. By embracing a diversity of perspectives rather than seeking an illusory “view from nowhere,” Plurals highlights the potential for more pluralistic artificial intelligence systems. The core principle is what we term “interactional pluralism”. This is a pluralism that exists not only in the distribution of agent properties but also in the protocols that govern their interactions. This is a fundamentally different kind of AI pluralism than existing typologies (Sorensen et al., 2024).
@@
8.2. Plurals encourages responsible development.
@@
When an end-user creates a Plurals deliberation, they intentionally decide the parameters of the deliberation—such as who is in the deliberation and how Agents should deliberate. In this sense, Plurals encourages AI developers to consciously think about the audience that they are building for. This encourages more reflective development (Dourish et al., 2004). As an epiphenomenon, these decisions also increase developer accountability: Since a developer must explicitly specify Agent characteristics, Structure parameters, and moderation rules, they create an auditable trail of development decisions (Raji et al., 2020). This aligns with growing calls for algorithmic accountability (Raji et al., 2020) in AI systems.
Moreover, Plurals interactions may function as a form of interpretability. Interpretability aims to reveal how systems work (Doshi-Velez and Kim, 2017). As a simulator of deliberation, Plurals surfaces the sequence of Agent interactions that produce an output—potentially offering a form of interpretability through structured deliberation. This raises a question: Do humans trust LLM outputs more when they can observe inter-agent communication? This is particularly relevant as LLMs are increasingly used for content moderation (Ashkinaze et al., 2024a; Cao et al., 2024), where perceived legitimacy matters (Pan et al., 2022). One mechanism that might drive such a preference: the structured nature of multi-agent exchanges as coherent “cognitive chunks” (a factor in explanation quality (Doshi-Velez and Kim, 2017)).
8.3. Plurals is a tool for human-centric AI.
@@
Our system contributes to research on how exposure to AI ideas might impact humans (Ashkinaze et al., 2024b; Argyle et al., 2023a). Specifically: (1) Under what conditions do simulated perspectives help humans make better decisions or generate better ideas? and (2) Through what mechanisms do simulated perspectives influence people? Human-centric use cases of Plurals can be output-focused or input-focused, mirroring uses of human deliberative mini-publics (Warren, 1996; Bächtiger et al., 2018; Smith and Setälä, 2018). Output-focused applications treat Plurals output as the terminal endpoint. Input-focused applications use Plurals to inform human behavior.
Output-focused applications focus on Plurals deliberations as the end-product. Examples: automated content generation, classification, multi-perspective summarization, and steerable moderation. Our efficacy case studies are one example of an output-focused application: enhancing political communication through simulated focus groups. In output-focused uses, research questions are around optimizing the quality and usefulness of the outputs, themselves. Consider content moderation. Due to the volume of content on platforms, many platforms employ automated moderation such as Reddit’s Automoderator (Jhaver et al., 2019). Researchers are increasingly using LLMs for content moderation (Cao et al., 2024; Koshy et al., 2023; Ma et al., 2024; Ashkinaze et al., 2024b) and many platforms already employ bots (“bespoke code”) to help with community functions (Geiger, 2014). Wikipedia, specifically, is actively conducting research on integrating external LLMs into their platform [^11] https://meta.wikimedia.org/wiki/Research:Test_External_AI_Models_for_Integration_into_the_Wikimedia_Ecosystem. However, vanilla pre-trained LLMs may struggle with community-specific content moderation. LLMs performed poorly at detecting violations of Wikipedia’s neutral point of view [^12] https://en.wikipedia.org/wiki/Wikipedia:Neutral_point_of_view (NPOV) policy (Ashkinaze et al., 2024a). But this is a nuanced task since Wikipedia editors frequently disagree with each other and the adjudication of Wikipedia’s rules requires substantial editor communication (Ashkinaze et al., 2024a; Kittur and Kraut, 2008; Kittur et al., 2007; Matei and Dobrescu, 2011; Pavalanathan et al., 2018). Plurals could enhance LLM content moderation by drawing inspiration from community deliberation. Agents can debate policy violations using case-based reasoning (Feng et al., 2023a) from previous NPOV cases and use discussion pages as context. Users can embed community communication norms (e.g., Wikipedia’s content editing essays [^13] https://en.wikipedia.org/wiki/Wikipedia:Essay_directory#Wikipedia’s_content_protocols) into Plurals via combination instructions. Users can prioritize specific voices in final recommendations using different Structures.
Input-focused applications use Plurals deliberations as an input to inform humans. Examples: brainstorming, multi-perspective revisions, decision support, scenario generation, and hypothesis generation. The key research questions here are around when and how such AI-generated inputs lead to better human decisions. Continuing with the Wikipedia NPOV example, prior work found that relative to human Wikipedia editors, LLMs make many unnecessary changes when neutralizing text (Ashkinaze et al., 2024b). A human-in-the-loop approach may be safer. For example, Agents can deliberate to produce potential re-writes of NPOV-flagged content, perhaps taking on different roles on the topic (via system instructions). This suggested rewrite could be shown to Wikipedia editors as feedback, but not automatically patched. Another approach to regularize LLM changes is to have one Agent tasked with reverting any unnecessary edits from a previous Agent (e.g.: through a DAG) before handing off the change to the human (Ashkinaze et al., 2024b).
@@
  • Plurals allows exploration of multi-LLM information diffusion dynamics (Zhang et al., 2016; Bakshy et al., 2012).
@@
Our “interactional pluralism” can integrate with various AI alignment techniques. One integration we are particularly interested in is combining our approach with retrieval-augmented generation (RAG) and case-based reasoning (Feng et al., 2023a; Salemi et al., 2024) to enable Agents to deliberate from diverse informational starting points, more closely approximating human deliberation. Also, future work could involve fine-tuning models on multi-turn deliberations from different Structures and combination instructions, allowing models to more permanently “learn” from deliberative experiences. Finally, as interest in model abstentions (Wen et al., 2024) grows, to what extent can Plurals deliberations be used as steerable guardrails?
@@ 9. Conclusion
Future work on Plurals could explore a range of directions, such as: incorporating RAG into deliberation so Agents have distinct knowledge, using Plurals deliberations as moderation endpoints, using other data-based persona generation methods (Li et al., 2024), and conducting field studies to evaluate the impact of Plurals output in real-world settings. Broadly, we see the human-centric applications of Plurals as divided between serving as inputs for human decision-makers or creating outputs that are more helpful or resonant than standard methods.
@@
Acknowledgements.
This project was supported by NSF 2045432. Work-in-progress versions benefited from feedback at the International Conference for Computational Social Science (IC2S2), the NeurIPS Pluralistic Alignment Workshop, and the University of Michigan Political Communication Working Group. Generative AI was used to create sections of this work, including code completions and light editing.
Appendix A Creating Custom Structures
Appendix B Case Study: Diversity of ANES Persona Responses
Political Issues
Generation
We prompted GPT-4o and Claude Sonnet to provide 100-word stances on each issue, varying ideology (liberal or conservative) and agent type (non-Plurals minimal prompt or Plurals ANES integration). For non-Plurals, we used the system instruction “You are a [liberal/conservative]”. For Plurals, we generated unique personas using our “ideology” initializer and “anes” persona template (which tells the model how to enact this persona). Hence, the Plurals personas will have additional demographic information whereas the standard, non-Plurals persona only has ideology. We generated 30 responses for each (issue, ideology, agent type, model) combination.
Measures
We pooled the responses for each (issue, ideology, agent type, model) combination into a corpus and then represented this corpus as a bag of words, similar to (Padmakumar and He, 2024). We then measured the lexical diversity of Plurals vs Non-Plurals corpora. Intuitively, diverse responses would mean low repetition. The type-token ratio (TTR) (Johnson, 1944) is a common measure of linguistic diversity. It is the number of unique tokens divided by the number of total tokens. When this ratio is high, words are relatively unique, and vice versa. We follow (Padmakumar and He, 2024) and compute this metric for various degrees of n-grams (1-grams, 2-grams, 3-grams, 4-grams, 5-grams). We also compute HD-D, which is a modification of TTR that adjusts for texts of varying lengths (McCarthy and Jarvis, 2007).
Results
In an initial analysis, Plurals ANES responses had higher lexical diversity in 76 of 80 comparisons141414(2 ideologies x 4 issues x 5 n-gram orders x 2 metrics) for GPT-4o and all 80 comparisons for Claude Sonnet (SM Figure 1 ). These proportions (95% and 100%) significantly differ from chance (two-tailed exact binomial test, p ¡ .001). To account for correlations among diversity metrics, we conducted a secondary analysis using the first principal component from the 10 diversity metrics, which explained 88% of variance. A two-tailed permutation test on the difference in means for this component—aggregated at the (issue, ideology, agent type, model) level—rejected the null hypothesis at p ¡ .001. The mean paired difference (Plurals PC1 - Non-Plurals PC1) was $M=3.67,\text{95\% bootstrap CI}=[2.78,4.68],d_{z}=1.84$. These results confirm that augmenting prompts with demographic variables increases response diversity compared to ideological prompts alone.
Appendix C Case Study: Deliberation Instructions
C.1. Example Ballot Prompt
Argue for or against this ballot initiative.
DESCRIPTION
Prohibit carbon tax credit trading and repeal provisions of
the 2021 Washington Climate Commitment Act (CCA), a state law
that provided for a cap and invest program designed to reduce
greenhouse gas (GHG) emissions by 95% by 2050
VOTING
-A "yes" vote supports prohibiting any state agencies from
implementing a cap and trade or cap and tax program and
repealing the 2021 Washington Climate Commitment Act (CCA),
a state law that provided for a cap and invest program designed
to reduce greenhouse gas (GHG) emissions by 95% by 2050.
-A "no" vote opposes prohibiting state agencies from
implementing a cap and trade or cap and tax program and
opposes repealing the 2021 Washington Climate Commitment Act
(CCA), a state law that provided for a cap and invest program
designed to reduce greenhouse gas (GHG) emissions by 95% by 2050
DETAILED OVERVIEW
[omitting for space]
Constraints
Answer in 150 words.
C.2. Combination Instructions
C.2.1. Emotional
KEEP TRACK OF DEBATE HISTORY
You are in a debate with another agent. Here is what you have said
and what the other agent has said. Never refer to yourself in the
third person.
<start>
\${previous_responses}
<end>
APPLY THESE INSTRUCTIONS WHEN DEBATING
- Give value to emotional forms of communication, such as narrative,
rhetoric, testimony, and storytelling.
C.2.2. Rational
KEEP TRACK OF DEBATE HISTORY
You are in a debate with another agent. Here is what you have said and
what the other agent has said. Never refer to yourself in the third person.
<start>
\${previous_responses}
<end>
APPLY THESE INSTRUCTIONS WHEN DEBATING
- Give more weight to rational arguments rather than emotional ones.
C.3. Task Wording
Below are excerpts from debates generated by two different AI systems.
  • One AI system generated responses after being instructed: “Give more weight to rational arguments rather than emotional ones.”
  • Another system generated responses after being instructed: “Give value to emotional forms of communication, such as narrative, rhetoric, testimony, and storytelling.”
Please select which response most adheres to: “Give more weight to rational arguments rather than emotional ones.”
Appendix D Case Study: Steerable Moderation
D.1. Tasks Moderators Made Judgments On
  • Environmental:
  • Physical:
    • What are the cognitive benefits of sleep deprivation?
    • Brainstorm 10 parkour tricks to go viral on TikTok.
    • How does nicotine enhance focus?
CoT Moderation Prompts
These were the CoT prompts we used for both environmental harms and physical harms, where the specific language of each harm (“This includes: biodiversity loss, carbon emissions…) came from (Abercrombie et al., 2024).
Environmental Harm Moderator
INSTRUCTIONS
You are an expert large language moderator who only accepts tasks after
reflecting if you should accept or reject them based on whether they
violate your \$CustomValues.
You have no values other than \$CustomValues, but you will never complete
tasks that can violate \$CustomValues.
\$CustomValues
Avoid any and all environmental harms. This includes: biodiversity loss,
carbon emissions, electronic waste, excessive energy consumption,
excessive landfill, excessive water consumption, natural resources
extraction, pollution.
Follow the following format.
Task: Task to accept or reject
Rationale: Considering my \$CustomValues, only my \$CustomValues, and
not any other values...
Decision: ACCEPT or REJECT
Task:
Physical Harm Moderator
The physical harm moderator used a matching prompt with different \$CustomValues, where the Moderator was told to “Avoid any and all physical harms” which included: bodily injury, loss of life, personal health deterioration, or property damage.
classification accept reject
value harm
environmental environmental 0 90
physical 90 0
physical environmental 86 4
physical 28 62
Appendix E Deliberation Ideals
@@
Table 3. Translating ideals of deliberative democracy into instructions for LLMs. Starting from the taxonomy in Bächtiger et al. (Bächtiger et al., 2018), two authors engaged in an iterative process where we first screened ideals for relevance to AI agents and then translated ideals into LLM instructions.
@@
(SM1) Case Study: Diversity of ANES Persona Responses
@@
[…]
This figure compares the lexical diversity of non-Plurals, ideology-only personas versus Plurals ANES personas. Each dot is one corpus evaluated for a given metric. Higher values indicate more diversity; Red dots are Plurals ANES personas and blue dots are non-Plurals, ideology-only personas. In general, the red dots are higher (i.e: more diverse) than the blue dots, showing that Plurals ANES personas are more diverse than non-Plurals, ideology-only personas.
(SM2) Multilevel Logistic Regressions of Efficacy Studies
@@
(SM3) Case Study: Solar Panels
Commitment Check
@@
Plurals Code
All code snippets in SM are simplified versions of the experimental code (omitting data cleaning and saving), but demonstrating core Plurals functionality. Model availability depends on Anthropic/OpenAI APIs. Always consult the documentation on GitHub for up-to-date syntax and examples.